OpenStudy (anonymous):

solve the equation : a^3(b-c)(x-b)(x-c) +b^3(c-a)(x-c)(x-a) + c^3(a-b)(x-a)(x-b)=0

OpenStudy (unklerhaukus):

\[a^3(b-c)(x-b)(x-c) +b^3(c-a)(x-c)(x-a) + c^3(a-b)(x-a)(x-b)=0\]

OpenStudy (unklerhaukus):

\[a^3\frac{(b-c)}{(x-a)} +b^3\frac{(c-a)}{(x-b)} + c^3\frac{(a-b)}{(x-c)}=0\]

OpenStudy (sriram):

really i am being lazy but x=a=b=c can be the answer

OpenStudy (sriram):

and to unkle rhaukus u cant divide an eq by smth which can be zero still u seem to be the only one trying respect

OpenStudy (unklerhaukus):

hmm, how are we going to fit the equation all on one line then?

OpenStudy (anonymous):

This equation might be zero on expanding.

OpenStudy (anonymous):

I am not really sure but

OpenStudy (anonymous):

-sigh- No :/ Wolfram doesn't get me Zero